# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ Appendix A Viscosity Solutions
This section reviews relevant definitions and results from the theory of viscosity solutions. General references for this material are [13, 7]. The main interest here is the initial-value problem
(57) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
@@
An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))\leq 0.$
A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))\geq 0.$
A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)\geq 0$ in the viscosity sense.
@@
An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (57) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
@@
A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (57) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
@@
Since the main theorem of this work, Theorem 1, involves problems with a discontinuous initial datum, a different definition of solution is needed. In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes. Specifically, for a function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,\infty)$ via the formulae
@@
A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (57) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)=(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)=(F_{\text{in}})_{*}$ in $\mathbb{R}$.
@@ A.1. Comparison Principle
Assume that the nonlinearity $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ is jointly continuous and nondecreasing in the second argument. If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
$\lim_{\delta\downarrow 0}\sup\left\{F(x,0)-G(y,0)\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\}\leq 0,$
then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (57) for some initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
The reader is referred to [16, Theorem 2.1], for instance, for the proof. The comparison principle implies that there is at most one continuous viscosity solutions of (57). Existence of the solution is also known in the present setting. However, there is no need to state an existence result here as it actually follows from the main convergence results (see the proof of Theorem 3).
Let $\overline{F}$ and $\underline{F}$ be the maximal sub- and minimal supersolution of (57):
@@
Notice that the comparison principle implies that $\overline{F}\leq\underline{F}$ and any viscosity solution (continuous or discontinuous) lies in between the two. This immediately implies a uniqueness criterion.
@@
Let $\overline{F}$ and $\underline{F}$ be the maximal subsolution and minimal supersolution of (57), respectively. If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (57) must also satisfy $F=\overline{F}$ in $\mathbb{R}\times(0,\infty)$. In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (57).
@@
As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\overline{F}\leq F\leq\underline{F}$ in $\mathbb{R}\times[0,\infty)$ in general. Therefore, in the present setting, $F=\overline{F}=\underline{F}$ for $t>0$, implying that there is at most one discontinuous viscosity solution. The fact that at least one exists is similarly immediate: The equality $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ actually implies that both $\overline{F}$ and $\underline{F}$ are discontinuous viscosity solutions. ∎
@@ A.2. Uniqueness for Nondecreasing Data
In this section, the PDE’s of interest here are shown to satisfy the uniqueness criterion of Proposition 9. For the Hamilton-Jacobi equations arising in the $\text{Beta}(2,1)$ case, the fact that almost-everywhere-continuous data leads to a unique solution at positive times is known from [10]. In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows from the characterization of nonnegative solutions of the porous medium equation. Short, simplified proofs are presented here for the reader’s convenience.
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\mathcal{F}(p,X)=-H(p)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying (i) and (ii) above, and if $\overline{F}$ and $\underline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (57), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
(60) $\overline{F}(x,t)=\underline{F}(x,t)=\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}.$
@@
When $F_{\text{in}}$ is uniformly Lipschitz continuous, the fact that the right-hand side of (60) determines the (continuous) viscosity solution is classical (see, e.g., [15, Section 10.3.4]). Otherwise, let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}})$ be a sequence of uniformly Lipschitz continuous functions such that
$F^{(N)}_{\text{in}}\leq F^{(N+1)}_{\text{in}}<F_{\text{in}}\quad\text{pointwise for each}\,\,N\in\mathbb{N},$
$\lim_{N\to\infty}F^{(N)}_{\text{in}}(x)=F_{\text{in}}(x)\quad\text{for each point of continuity}\,\,x\,\,\text{of}\,\,F_{\text{in}}.$
If $\{F_{N}\}$ are the corresponding solutions of (57), then the comparison principle implies that $F_{N}\leq F_{N+1}\leq\overline{F}$ for each $N$. By the stability of viscosity solutions, the limit $F=\lim_{N\to\infty}F_{N}$ satisfies $\partial_{t}F-H(\partial_{x}F)=0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
At the same time, recall that the superlinearity of $H$ implies that $L$ is also superlinear. Therefore, for any $(x,t)$, it is possible to write $F_{N}(x,t)=F^{(N)}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$. If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
$F_{\text{in}}(y-\delta)=\lim_{N\to\infty}F_{N}(y-\delta)\leq\liminf_{N\to\infty}F^{(N)}_{\text{in}}(y_{N})\quad\text{for almost every}\,\,\delta>0,$
from which it follows that
$\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}\leq\lim_{N\to\infty}F_{N}(x,t)=F(x,t)\leq\overline{F}(x,t).$
$\underline{F}$ can be treated similarly: Let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}}$ be uniformly Lipschitz functions such that $F_{\text{in}}<F^{(N+1)}_{\text{in}}\leq F^{(N)}_{\text{in}}$ for each $N$ and $F_{\text{in}}(x)=\lim_{N\to\infty}F^{(N)}_{\text{in}}(x)$ at points of continuity $x$ of $F_{\text{in}}$. This time, if $F_{N}$ is the corresponding solution of (57), then, for any point of continuity $y$ of $F_{\text{in}}$ and any $(x,t)\in\mathbb{R}\times(0,\infty)$,
$\underline{F}(x,t)\leq\lim_{N\to\infty}F_{N}(x,t)\leq\lim_{N\to\infty}F_{\text{in}}^{(N)}(y)+tL(\frac{x-y}{t})=F_{\text{in}}(y)+tL(\frac{x-y}{t}).$
Since $F$ is right-continuous and points of continuity are dense, this proves $\underline{F}$ is bounded above by the right-hand side of (60). ∎
Next, consider the case of the nonlinearity
(61) $\mathcal{F}(p,X)=a|p|^{m-1}X$
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\mathcal{F}$ has the form (61) for some $m>1$, then the maximal subsolution $\overline{F}$ and minimal supersolution $\underline{F}$ of (57) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
$\overline{F}(x,t)=\underline{F}(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy,$
@@
When $F_{\text{in}}$ is uniformly Lipschitz, this result is a special case of [3, Theorem 3.1].
Unless otherwise noted, basic properties of the PME used in the proof that follows can be found in the monograph [21].
@@
First, assume that $F_{\text{in}}$ is smooth and $\partial_{x}F_{\text{in}}$ is bounded and nowhere vanishing. In this case, classical results on the PME imply that there is a unique smooth solution $\rho$ of (64), which is positive and satisfies $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for all $t>0$. Since $\rho$ is smooth, the function $F(x,t)=\int_{-\infty}^{x}\rho(y,t)\,dy$ is also, and, thus, it satisfies $\partial_{t}F-a|\partial_{x}F|^{m-1}\partial_{x}^{2}F=\partial_{t}F-\frac{a}{m}\partial_{x}(\partial_{x}F)^{m}=0$ classically. As is well-known, being a classical solution, it is also a viscosity solution.

### v2

@@ Appendix A Viscosity Solutions
This section reviews relevant definitions and results from the theory of viscosity solutions. General references for this material are [15, 10]. The interest here is in the initial-value problem
(61) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x),\end{array}\right$.
where $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and the diffusion term $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and the reaction term $Q:\mathbb{R}\to\mathbb{R}$ are continuous functions, which satisfy the assumptions, for some $L\geq 0$,
(62) $\mathcal{F}(v,w_{1})\leq\mathcal{F}(v,w_{2})\quad\text{for each}\,\,v,w_{1},w_{2}\in\mathbb{R}\,\,\text{with}\,\,w_{1}<w_{2},$
(63) $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|\,\,\text{for each}\,\,q,q^{\prime}\in\mathbb{R},$
(64) $\mathcal{F}(0,0)=0,\quad Q(0)\geq 0,\quad\text{and}\quad Q(1)\leq 0.$
@@
An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from above (that is, $F(x,t)\leq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\leq 0.$
A lower semi-continuous function $G:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential inequality $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial_{x}^{2}G)+Q(G)\geq 0$ in the viscosity sense if it has the following property: If $\varphi$ is a smooth function defined in some open set $U$ and if there is a point $(x_{0},t_{0})\in U$ at which $\varphi$ touches $F$ from below (that is, $F(x,t)\geq\varphi(x,t)$ for all $(x,t)\in U$ with equality when $(x,t)=(x_{0},t_{0})$), then
$\partial_{t}\varphi(x_{0},t_{0})-\mathcal{F}(\partial_{x}\varphi(x_{0},t_{0}),\partial_{x}^{2}\varphi(x_{0},t_{0}))+Q(\varphi(x_{0},t_{0}))\geq 0.$
A continuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ is said to satisfy the differential equation $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0$ in the viscosity sense if it satisfies both $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\geq 0$ in the viscosity sense.
@@
An upper semicontinuous function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ said to be a viscosity subsolution of (61) if it satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
@@
A lower semicontinuous function $G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a viscosity supersolution of (61) if it satisfies $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and
@@
Since the main results of this work, Theorems 1 and 2, involve problems with a discontinuous initial datum, a different definition of solution is needed. In order to state it, it is necessary to recall the notions of upper and lower semicontinuous envelopes. Specifically, for a function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$, the upper and lower semicontinuous envelopes $F^{*}$ and $F_{*}$ are the functions defined in $\mathbb{R}\times[0,\infty)$ via the formulae
@@
A function $F:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ is said to be a discontinuous viscosity solution of (61) if $\partial_{t}F^{*}-\mathcal{F}(\partial_{x}F^{*},\partial_{x}^{2}F^{*})+Q(F^{*})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; $\partial_{t}F_{*}-\mathcal{F}(\partial_{x}F_{*},\partial_{x}^{2}F_{*})+Q(F_{*})\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$; and $F^{*}(\cdot,0)\leq(F_{\text{in}})^{*}$ and $F_{*}(\cdot,0)\geq(F_{\text{in}})_{*}$ in $\mathbb{R}$.
@@ A.1. Comparison Principle
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63). If $F,G:\mathbb{R}\times[0,\infty)\to\mathbb{R}$ are respectively a bounded upper and a bounded lower semicontinuous function such that $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial_{x}^{2}F)+Q(F)\leq 0$ and $\partial_{t}G-\mathcal{F}(\partial_{x}G,\partial^{2}_{x}G)+Q(G)\geq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$ and if $M$ is the constant defined by
$M=\lim_{\delta\downarrow 0}\sup\left\{(F(x,0)-G(y,0))_{+}\,\mid\,x,y\in\mathbb{R},\,\,|x-y|\leq\delta\right\},$
then, for any $t>0$,
(66) $\sup\left\{F(x,t)-G(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq Me^{Lt}$
In particular, if $F$ and $G$ are respectively bounded viscosity sub- and supersolutions of (61) for some fixed initial datum $F_{\text{in}}$, then $F\leq G$ pointwise in $\mathbb{R}\times[0,\infty)$.
Proof.
The reader can find a proof of this version of the theorem with $L=0$, $M=0$ in [19]. The extension to the case $L>0$, $M=0$ is routine (see also [10, Section 5.2]). The case $L>0$, $M>0$ can be reduced to the case $L>0$, $M=0$ using an analogue of the argument appearing in the proof of Lemma 5 below, namely, using that the function $\tilde{F}=F-Me^{Lt}$ satisfies $\partial_{t}\tilde{F}-\mathcal{F}(\partial_{x}\tilde{F},\partial_{x}^{2}\tilde{F})+Q(\tilde{F})\leq 0$ in $\mathbb{R}\times(0,+\infty)$. ∎
The comparison principle implies, in particular, that there is at most one viscosity solution of (61) if $F_{\text{in}}$ is continuous. It is known that, indeed, a solution exists in this case, and the assumption (64) ensures that, for each $t>0$, $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$.
Proposition 12.
Assume that $\mathcal{F}$ and $Q$ satisfy (62), (63), and (64). For any continuous $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, there is a unique viscosity solution $F$ of (61). Further, for any $t>0$, the function $F(\cdot,t)$ is in $CDF(\overline{\mathbb{R}})$.
Proof.
Existence of a solution taking values in $[0,1]$ follows from Perron’s Method, which is applicable since the constant functions 0 and $1$ are, respectively, viscosity sub- and supersolutions by (64). By the comparison principle, the solution is unique and it takes values in $[0,1]$. Finally, for any $y\in\mathbb{R}$, the function $(x,t)\mapsto F(x+y,t)$ is a viscosity solution of (61) with the initial datum $x\mapsto F_{\text{in}}(x+y)$. Thus, since $F_{\text{in}}$ is nondecreasing, by the comparison principle, $F(x+y,t)\leq F(x,t)$ for all $t\geq 0$ if $y\geq 0$. Since $F$ is continuous, this proves $F(\cdot,t)\in CDF(\overline{\mathbb{R}})$ for each $t>0$. ∎
In general, let $\underline{F}$ and $\overline{F}$ be the maximal sub- and minimal supersolution of (61):
@@
Notice that the comparison principle implies that $\underline{F}\leq\overline{F}$ and any viscosity solution (continuous or discontinuous) lies in between the two. This immediately implies a uniqueness criterion.
@@
Assume that $\mathcal{F}$ and $Q$ satisfy assumptions (62) and (63). Let $\underline{F}$ and $\overline{F}$ be the maximal subsolution and minimal supersolution of (61), respectively, for some given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,\infty)$, then any discontinuous viscosity solution $F$ of (61) must also satisfy $F=\underline{F}$ in $\mathbb{R}\times(0,\infty)$. In particular, in this case, modulo the freedom to choose the value at discontinuities of $F_{\text{in}}$ at $t=0$, there is a unique discontinuous viscosity solution of (61), which is continuous in $\mathbb{R}\times(0,+\infty)$.
@@
As in the discussion preceding the statement, if $F$ is any discontinuous viscosity solution, then $\underline{F}\leq F\leq\overline{F}$ in $\mathbb{R}\times[0,\infty)$ in general. Therefore, in the present setting, $F=\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, implying that there is at most one discontinuous viscosity solution (up to arbitrary choices at $t=0$ at points where $F_{\text{in}}$ is discontinuous).
The fact that at least one exists is almost immediate: Since $\underline{F}$ is a supremum of subsolutions, the upper semicontinuous envelope $\underline{F}^{*}$ satisfies $\partial_{t}\underline{F}^{*}-\mathcal{F}(\partial_{x}\underline{F}^{*},\partial_{x}^{2}\underline{F}^{*})+Q(\underline{F}^{*})\leq 0$ in $\mathbb{R}\times(0,+\infty)$, and solutions of (61) with initial data larger than $F_{\text{in}}$ can be used to prove $\underline{F}^{*}(\cdot,0)=F_{\text{in}}$ at $t=0$. Therefore, by comparison, $\underline{F}^{*}\leq\overline{F}$ in $\mathbb{R}\times[0,+\infty)$. A similar argument shows that $\partial_{t}\overline{F}_{*}-\mathcal{F}(\partial_{x}\overline{F}_{*},\partial^{2}_{x}\overline{F}_{*})+Q(\overline{F}_{*})\geq 0$ in $\mathbb{R}\times(0,+\infty)$ and $\overline{F}_{*}(\cdot,0)$ equals the lower semicontinuous envelope of $F_{\text{in}}$, so $\overline{F}_{*}\geq\underline{F}$. Since $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$ and the upper (resp. lower) semicontinuous envelope of a function is always larger (resp. smaller) than the function itself, this proves $\underline{F}$ is continuous in $\mathbb{R}\times(0,+\infty)$ and a discontinuous viscosity solution of (61). ∎
The following lemma will be used below to verify the identity $\underline{F}=\overline{F}$.
Lemma 1.
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. Suppose that $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ are two sequences of continuous functions in $CDF(\overline{\mathbb{R}})$ such that
(67) $F^{(N-1),-}_{\text{in}}\leq F^{(N),-}_{\text{in}}\leq F_{\text{in}}\leq F^{(N),+}_{\text{in}}\leq F^{(N-1),+}_{\text{in}}\quad\text{pointwise for each}\,\,N,$
(68) $\lim_{N\to+\infty}F^{(N),+}_{\text{in}}=\lim_{N\to+\infty}F^{(N),-}_{\text{in}}=F_{\text{in}}\quad\text{vaguely.}$
Let $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ be the (continuous) viscosity solutions of (61) with respective initial data $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$. If the limit of the two sequences coincides, that is, if
$F:=\lim_{N\to+\infty}F^{(N),+}=\lim_{N\to+\infty}F^{(N),-}\quad\text{pointwise in}\,\,\mathbb{R}\times(0,+\infty),$
then the maximal subsolution and minimal supersolution of (61) with initial datum $F_{\text{in}}$ are equal in $\mathbb{R}\times(0,+\infty)$ and, in particular, they both equal the limit $F$.
Proof.
This is immediate from the fact that $F^{(N),-}\leq\underline{F}\leq\overline{F}\leq F^{(N),+}$ for any $N$ by comparison. ∎
@@ A.2. Uniqueness for Nondecreasing Data
In this section, the PDE’s of interest here are shown to satisfy the uniqueness criterion of Proposition 13. For the Hamilton-Jacobi equations arising in the $\text{Beta}(2,1)$ case, the fact that almost-everywhere-continuous data leads to a unique solution at positive times is known from [11]. In the case of the second-order equation in the $\text{Beta}(2,2)$ case, the uniqueness of solutions follows, for instance, from the characterization of nonnegative solutions of the porous medium equation. Short, simplified proofs are presented here for the reader’s convenience.
Case: $Q=0$. The discussion begins with the classical case when the reaction term $Q=0$.
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$. If $\mathcal{F}(v,w)=-H(v)$ for a convex $H:\mathbb{R}\to\mathbb{R}$ satisfying $H(0)=0$ and (69), and if $\underline{F}$ and $\overline{F}$ are the maximal subsolution and minimal supersolution, respectively, of (61), then, for any $(x,t)\in\mathbb{R}\times(0,\infty)$,
(70) $\underline{F}(x,t)=\overline{F}(x,t)=\inf\left\{F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)\,\mid\,y\in\mathbb{R}\right\}.$
@@
Furthermore, the following statements hold:
  • (i)
    For any $t>0$, the functions $\underline{F}(\cdot,t)$ and $F_{\text{in}}$ coincide on the set $\{-\infty,+\infty\}$.
  • (ii)
    If $\{F^{(n)}_{\text{in}}\}$ is a sequence in $CDF(\overline{\mathbb{R}})$ converging vaguely to $F_{\text{in}}$, and if $\{\underline{F}^{(n)}\}$ are the corresponding maximal subsolutions of (61), then, for any compact set $K\subseteq\mathbb{R}\times(0,+\infty)$,
    $\lim_{n\to\infty}\sup\left\{|\underline{F}^{(n)}(x,t)-\underline{F}^{(n)}(x,t)|\,\mid\,(x,t)\in K\right\}=0.$
@@
Let $F$ be the function defined in the righthand side of (70). When $F_{\text{in}}$ is continuous, the fact that $F$ is the (continuous) viscosity solution is classical (see, e.g., [18, Section 10.3.4]). Otherwise, let $\{F^{(N),+}_{\text{in}}\}$ and $\{F_{\text{in}}^{(N),-}\}$ be two sequences converging to $F_{\text{in}}$ as in Lemma 1, and let $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ be the corresponding solutions given by (70). For any point of continuity $y$ of $F_{\text{in}}$, any $x\in\mathbb{R}$, and any $t>0$,
$F_{\text{in}}(y)+tL\left(\frac{x-y}{t}\right)=\lim_{N\to\infty}F_{\text{in}}^{(N),-}(y)+tL\left(\frac{x-y}{t}\right)\geq\lim_{N\to\infty}F^{(N),+}(x,t).$
Therefore, since $F_{\text{in}}$ is right-continuous, this proves $\lim_{N\to\infty}F^{(N),+}\leq F$ pointwise in $\mathbb{R}\times(0,+\infty)$. Next, by superlinearity of $L$, for any $(x,t)$, it is possible to write $F^{(N),-}(x,t)=F^{(N),-}_{\text{in}}(y_{N})+tL((x-y_{N})/t)$ for some bounded sequence $\{y_{N}\}$. If $y_{N}\to y$ as $N\to\infty$ (say, along a subsequence), then
$F_{\text{in}}(y-\delta)=\lim_{N\to\infty}F^{(N),-}(y-\delta)\leq\liminf_{N\to\infty}F^{(N),-}_{\text{in}}(y_{N})\quad\text{for almost every}\,\,\delta>0,$
from which it follows that $F(x,t)\leq\lim_{N\to+\infty}F^{(N),-}(x,t)$. Therefore, by Lemma 1, $F=\underline{F}=\overline{F}$.
Item (i) follows directly from the representation (70), the superlinearity of $L$, and the fact that $0\leq F_{\text{in}}\leq 1$. Item (ii) can be proved using the comparison principle, Theorem 4, in conjunction with the half-relaxed limit method; the interested reader can compare with the proof of Corollary 4 below. ∎
Next, let $m>1$ and consider the case of the nonlinearity
(71) $\mathcal{F}(v,w)=a|v|^{m-1}w$
@@
Let $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ and assume $Q=0$. If $\mathcal{F}$ has the form (71) for some $m>1$, then the maximal subsolution $\underline{F}$ and minimal supersolution $\overline{F}$ of (61) are given for $(x,t)\in\mathbb{R}\times(0,\infty)$ by the formula
(72) $\underline{F}(x,t)=\overline{F}(x,t)=F_{\text{in}}(-\infty)+\int_{-\infty}^{x}\rho(y,t)\,dy,$
@@
Furthermore, the following statements hold:
  • (i)
    For any $t>0$, the functions $\underline{F}(\cdot,t)$ and $F_{\text{in}}$ coincide on the set $\{-\infty,+\infty\}$.
  • (ii)
    If $\{F_{\text{in}}^{(n)}\}$ is some sequence in $CDF(\overline{\mathbb{R}})$ converging vaguely to $F_{\text{in}}$, and if $\{\underline{F}^{(n)}\}$ are the corresponding maximal subsolutions of (61), then, for any compact set $K\subseteq\mathbb{R}\times(0,+\infty)$,
    $\lim_{n\to+\infty}\sup\left\{|\underline{F}^{(n)}(x,t)-\underline{F}(x,t)|\,\mid\,(x,t)\in K\right\}=0.$
When $F_{\text{in}}$ is uniformly Lipschitz, the representation formula obtained above is a special case of [3, Theorem 3.1].
@@
The uniqueness and claimed representation formula for the solution follows from the classical uniqueness theorem for the PME by Dahlberg and Kenig [16]. To see this, first, note that if $\partial_{x}F_{\text{in}}$ is smooth and positive, then it is well-known that the solution $\rho$ of the PME is smooth (up to time $t=0$), hence the antiderivative given by (72) is a smooth solution of (61). It is classical that smooth solutions are viscosity solutions. It is also classical in this setting that $\int_{-\infty}^{\infty}\rho(y,t)\,dy=F_{\text{in}}(+\infty)-F_{\text{in}}(-\infty)$ for each $t>0$, as in item (i).
It will be convenient later in the proof to note that, in this one-dimensional setting, since the initial data is a subprobability measure, the pressure $\rho^{m-1}$ associated with any of the smooth solutions above satisfies the following estimates
(76) $\tau^{\frac{m-1}{m+1}}|\rho^{m-1}|+\tau^{\frac{m}{m+1}}|\partial_{x}(\rho^{m-1})|+\tau^{\frac{2m}{m+1}}|\partial_{t}(\rho^{m-1})|\leq C\quad\text{in}\,\,\mathbb{R}\times(\tau,+\infty),$
where $C>0$ is a universal constant and $\tau>0$ is arbitrary (see [24, Chapter 15]).